# When Does Multi-Task Learning Fail? Quantifying Data Imbalance and Task Independence in Metal Alloy Property Prediction

### v2

@@ 5.2.2 Different Optimization Landscapes
Classification with cross-entropy loss has different optimization properties than regression with MSE loss. The classification loss is more robust to the gradient interference caused by multi-task optimization, allowing the model to find solutions that perform well on classification despite suboptimal regression performance.
@@ 5.3.3 Recommendation 3: Validate MTL Assumptions Before Deployment
5.4 Toward a Materials Taskonomy
Our findings suggest the need for a systematic “taskonomy” of material properties, analogous to that developed by Zamir et al. [zamir2018taskonomy] for computer vision tasks. Rather than assuming all material properties are related because they derive from atomic-level physics, we should empirically map which properties share sufficient representational structure to benefit from joint learning.
Based on our analysis and the underlying physics, we propose organizing alloy properties into clusters by physical mechanism:
@@ 5.4 Toward a Materials Taskonomy
Key insight: The negative transfer observed in our study occurred precisely because we attempted to bridge the Electronic cluster (resistivity) and the Mechanical cluster (hardness) using a shared representation dominated by Electronic-task data. Properties within the same cluster—such as hardness and yield strength—share common functional dependencies and are likely to exhibit positive transfer. Properties across clusters follow different mathematical forms and may actively interfere when forced to share representations.
@@
This taskonomy framework has practical implications: when constructing MTL models for materials discovery, practitioners should preferentially group properties within the same physical cluster. Cross-cluster MTL should be approached with caution and requires explicit validation against independent baselines.
@@ 5.5.1 Dataset Limitations
Our analysis is limited to one dataset (AI-Hub Korea metal alloys) with one specific set of properties. Different property combinations or different material systems may exhibit stronger inter-task relationships. Future work should validate our findings across diverse materials datasets.
@@ 5.5.2 Effectiveness of DIR Mitigation Techniques
PCGrad [yu2020gradient]: Highly effective for minority-task protection. The +12.4% improvement on hardness confirms that gradient conflicts were the primary bottleneck for this task. However, it comes at a cost to majority-task performance (-2.5% on resistivity).
@@
GradNorm [chen2018gradnorm]: Provides balanced improvements across all tasks through dynamic weight adjustment. Combined with LDS, achieves the best overall performance profile.
@@
LDS and Balanced MSE [yang2021delving, ren2022balanced]: Modest improvements for classification (+5.3% F1) but limited benefit for regression tasks when used alone.
@@ 6 Conclusion
This study critically examined the effectiveness of multi-task learning for metal alloy property prediction, challenging the common assumption that MTL universally improves performance when properties share underlying physics. By integrating our empirical findings with the Deep Imbalanced Regression (DIR) framework [yang2021delving], we provide both practical guidelines and theoretical insight into the mechanisms of negative transfer.
Using a comprehensive dataset of 54,028 metal alloy samples with three target properties (electrical resistivity, Vickers hardness, and amorphous-forming ability), we compared independent single-task models, standard MTL with hard parameter sharing, and structured MTL with learned task relation graphs. Our key findings are:
@@
MTL significantly degrades regression performance. For resistivity prediction, dropped from 0.897 to 0.844 (). For hardness, the degradation was more severe: dropped from 0.832 to 0.694 ().
@@
MTL significantly improves classification performance. For amorphous-forming ability prediction, F1 score improved from 0.703 to 0.744 (), with recall improving by 17% ().
@@
Learned task relations reveal property independence. The structured MTL model learned near-zero inter-task weights (0.006), indicating that these alloy properties do not share sufficient common structure to benefit from joint learning.
@@
Data imbalance is a critical factor. The 65:1 ratio between resistivity and hardness samples causes the dominant task to overwhelm minority tasks during optimization, contributing to negative transfer.
@@
DIR mitigation techniques effectively address negative transfer. PCGrad improved hardness by 12.4% (0.761 0.855), confirming that gradient conflicts were the primary bottleneck. LDS+GradNorm achieved the best overall balance with improvements across all tasks. However, combining techniques does not guarantee additive benefits—LDS+PCGrad showed negative synergy.
These findings provide important guidelines for practitioners in materials informatics:
1.
Use independent models when precise property prediction is required
2.
Use MTL for classification/screening tasks where improved recall is valuable
3.
Validate MTL assumptions empirically rather than assuming positive transfer
4.
Consider data imbalance as a primary factor in MTL effectiveness
5.
Apply MTL within property clusters (electronic, mechanical, thermodynamic) rather than across them
@@
Use PCGrad when maximizing minority-task performance is the priority
7.
Use LDS+GradNorm when balanced multi-task performance is desired
8.
@@
We propose a “materials taskonomy” framework that organizes properties by their underlying physics. Properties within the same cluster—such as hardness and yield strength (mechanical) or resistivity and thermal conductivity (electronic)—are expected to benefit from joint learning, while cross-cluster MTL requires careful validation.
Our evaluation of DIR techniques—including LDS, Balanced MSE, PCGrad, and GradNorm—demonstrates that gradient conflict resolution can substantially mitigate negative transfer, with PCGrad achieving +12.4% improvement on the minority task. Future work should explore active learning approaches [rao2022machine] that curate informative training distributions, and sequential transfer learning [kamnis2024transformer] where models are pre-trained on large datasets and fine-tuned for specific properties.
Our work contributes to a more nuanced understanding of multi-task learning in materials science, moving beyond the assumption that “more data is always better” to recognize that how tasks are combined matters as much as the data itself. By providing both positive and negative results with rigorous statistical analysis, we hope to help researchers make more informed decisions about when to apply MTL in their materials discovery pipelines.
@@
The metal alloy property datasets used in this study are available from the AI-Hub Korea platform [aihub2022material]. Access requires free registration. The datasets include: (1) electrical resistivity measurements (52,388 samples), (2) Vickers hardness measurements (800 samples), and (3) amorphous-forming ability labels (840 samples). All experimental code, including implementations of LDS, FDS, PCGrad, GradNorm, and soft parameter sharing, is available at https://github.com/[repository] under the MIT license.
CRediT Author Statement
Author One: Conceptualization, Methodology, Software, Writing - Original Draft. Author Two: Validation, Investigation, Writing - Review & Editing. Author Three: Supervision, Project Administration, Writing & Editing.
@@
Acknowledgments
This research was supported by [Funding Agency]. The authors thank [acknowledgments].

### v3

@@ 5.2.2 Different Optimization Landscapes
Classification with cross-entropy loss exhibits fundamentally different optimization dynamics compared to regression with MSE. Although we applied the same aggressive inverse weighting () to both minority tasks, this strategy proved effective for classification while detrimental to regression. The cross-entropy loss, defined over bounded probabilities, scales logarithmically; crucially, its gradients are naturally bounded by the difference between predicted probabilities and labels. This makes it inherently robust to large scalar multipliers. In contrast, MSE is unbounded and quadratic, where gradients scale linearly with the error magnitude. Consequently, heavy weighting in regression amplifies these errors, leading to gradient explosion. Thus, for the amorphous task, the heavy weighting successfully mitigated data imbalance without the optimization instability observed in the hardness task.
@@ 5.3.3 Recommendation 3: Validate MTL Assumptions Before Deployment
5.4 A Hypothesis for Materials Property Clustering
Our findings motivate—but do not validate—a hypothesis about organizing material properties for multi-task learning, inspired by the “taskonomy” developed by Zamir et al. [28] for computer vision tasks.
Hypothesis: Alloy properties can be organized by underlying physical mechanism, with within-mechanism properties (e.g., hardness and yield strength) expected to exhibit positive MTL transfer, while cross-mechanism properties (e.g., resistivity and hardness) are expected to exhibit negative transfer.
Evidence from current work: The three properties we studied—resistivity (electronic), hardness (mechanical), and amorphous-forming ability (thermodynamic)—span different physical mechanisms. All exhibited negative transfer or task independence, consistent with this hypothesis.
Validation status: This hypothesis remains unvalidated. Rigorous testing would require systematic experiments on within-mechanism property pairs, which our current dataset does not support. We present this as a direction for future investigation rather than as an established framework.
Based on the underlying physics, we tentatively propose organizing alloy properties into clusters by physical mechanism:
@@ 5.4 A Hypothesis for Materials Property Clustering
Interpretation: The negative transfer observed in our study is consistent with crossing cluster boundaries: we combined the Electronic cluster (resistivity) and the Mechanical cluster (hardness) using a shared representation dominated by Electronic-task data. If our hypothesis holds, properties within the same cluster—such as hardness and yield strength—would share common functional dependencies and exhibit positive transfer. Properties across clusters would follow different mathematical forms and may interfere when forced to share representations.
@@
Practical implications (pending validation): If validated, this clustering hypothesis would suggest that practitioners constructing MTL models for materials discovery should preferentially group properties within the same physical cluster. Cross-cluster MTL would require explicit validation against independent baselines. However, we emphasize that these implications are speculative until the hypothesis is tested on appropriate datasets.
5.4.1 The Pitfall of Derived Properties
To empirically test the property clustering hypothesis, we derived thermal conductivity from resistivity using the Wiedemann-Franz law:
where W/K2 is the Lorenz number and K. This creates a within-mechanism property pair (both governed by electronic transport) to compare against the cross-mechanism baseline.
Experimental Design:
1.
Scenario A (Cross-mechanism): Resistivity + Hardness + Amorphous
2.
Scenario B (Within-mechanism): Resistivity + Thermal Conductivity (derived) + Amorphous
We trained MTL models for both scenarios using 5 random seeds with 200 epochs and early stopping (patience=30). Table 13 shows the results:
| Scenario | Configuration | Resistivity | Transfer Type |
| A (Cross-mech) | Res+Hard+Amor | 0.897 0.006 | Cross-mechanism |
| B (Within-mech) | Res+Therm+Amor | 0.874 0.013 | Within-mechanism |
| (B - A) | – | 0.023 | – |
Key Finding: The within-mechanism configuration (Scenario B) performed worse than the cross-mechanism configuration (Scenario A) by 2.3% . This contradicts the naive property clustering hypothesis.
Explanation: Information Redundancy. The derived thermal conductivity is a direct mathematical transformation of resistivity (). Consequently:
1.
The model learns the same underlying composition-to-property mapping twice
2.
Gradients from the auxiliary task are linear transformations of the primary task gradients
3.
No new information is provided to improve the shared representation
In contrast, hardness—despite being from a different physical mechanism—provides orthogonal information that captures mechanical/structural aspects of alloys not encoded in resistivity.
Practical Implication: This result serves as a warning to practitioners: augmenting MTL datasets using known physical laws (e.g., Wiedemann-Franz, Hall-Petch) does not improve performance and may degrade it. For positive transfer, auxiliary tasks must provide informationally independent supervision, not mathematically derived quantities.
Refined Hypothesis: We propose that successful MTL for materials properties requires not just shared physical mechanisms, but informationally independent measurements. Future work should test this refined hypothesis using independently measured (not derived) within-mechanism property pairs.
5.5 Comparison with Recent Contradictory Work
Our findings of negative transfer for regression tasks contrast with several recent publications reporting positive MTL outcomes in materials science. A 2025 study in Acta Materialia [21] demonstrated substantial improvements (+37.5%) using MTL for superalloy property prediction across six thermodynamic and microstructural properties. Similarly, Debnath and Reinhart [6] reported modest but consistent improvements when applying MTL to high-entropy alloy yield strength and elongation prediction.
These contrasting results are not necessarily contradictory; rather, they highlight the importance of understanding when MTL succeeds versus fails. We identify three key factors that may explain the divergent outcomes:
1.
Property selection: The superalloy study predicted properties within the same physical domain (thermodynamic/microstructural properties), whereas we combined properties from different domains (electronic, mechanical, phase stability). This is consistent with our property clustering hypothesis: within-domain properties may share sufficient representational structure to benefit from joint learning, while cross-domain properties may not.
2.
Data balance: Different sample size ratios across tasks determine the severity of negative transfer. Our extreme 65:1 imbalance may have created conditions particularly unfavorable for MTL. Studies with more balanced task sizes may avoid this failure mode.
3.
Material system: Property relationships may be composition-dependent. Some alloy families (e.g., Ni-based superalloys) may exhibit stronger inter-property correlations due to their specific chemistry and physics, making them more amenable to MTL.
These considerations suggest that MTL success depends critically on property selection and data characteristics, not just the materials domain. Our negative results do not invalidate MTL for materials science broadly; rather, they identify conditions under which caution is warranted. A complete understanding will require systematic studies across diverse material systems and property combinations.
@@ 5.6.1 Dataset Limitations
Our analysis is limited to one dataset (AI-Hub Korea metal alloys) with three specific properties (resistivity, hardness, amorphous-forming ability). While limiting the study to a single experimental database restricts generalizability, it was a necessary constraint to perform the functional form analysis in Section 3. Large-scale computational databases (e.g., OQMD, Materials Project) were excluded from this study because they lack the specific experimental noise and measurement-availability imbalance (65:1) that characterize real-world experimental archives. Future work should seek to replicate these findings on independent experimental alloy databases (e.g., Citrination) rather than computational proxies.
This scope constrains the generalizability of our conclusions in several important ways:
1.
Single dataset: We cannot claim that our findings represent general principles about MTL in materials science. Different material systems, composition spaces, and property combinations may yield different outcomes.
2.
Limited property coverage: With only three properties spanning different physical domains, we cannot fully test whether within-domain MTL (e.g., hardness + yield strength) would succeed where our cross-domain approach failed. Our preliminary test using derived thermal conductivity (Section 5.4.1) suggests that independently measured within-mechanism properties are needed for valid hypothesis testing.
3.
Specific imbalance characteristics: Our extreme 65:1 data imbalance may represent a particularly challenging scenario. Datasets with more balanced task sizes may not experience the same negative transfer.
4.
Contradictory recent evidence: As discussed above, recent studies on different material systems report opposite results (MTL improvements rather than degradation), underscoring that our findings may be dataset- and property-specific.
We explicitly acknowledge that our results should be interpreted as observations on a specific dataset rather than as universal principles. Validation across diverse material systems is essential before general guidelines can be established.
@@ 5.6.2 Effectiveness of DIR Mitigation Techniques
PCGrad [27]: Highly effective for minority-task protection. The +12.4% improvement on hardness confirms that gradient conflicts were the primary bottleneck for this task. However, it comes at a cost to majority-task performance (-2.5% on resistivity).
@@
GradNorm [4]: Provides balanced improvements across all tasks through dynamic weight adjustment. Combined with LDS, achieves the best overall performance profile.
@@ 5.6.5 Feature Distribution Smoothing
5.6.6 Testing with Independent Measurements
Our derived property experiment (Section 5.4.1) highlights the critical distinction between mechanistic similarity and informational independence. The failure of within-mechanism MTL using derived thermal conductivity demonstrates that shared physics alone is insufficient—auxiliary tasks must provide genuinely independent learning signals. Future work should:
1.
Collect independently measured within-mechanism property pairs (e.g., resistivity and Seebeck coefficient, both from electronic transport but measured separately)
2.
Test whether informational independence, rather than physical mechanism, is the true predictor of MTL transfer success
3.
Develop metrics to quantify the “information overlap” between property datasets prior to MTL training
4.
Investigate whether the negative result from derived properties generalizes to other physical laws (e.g., Hall-Petch relation for grain size and yield strength)
@@ 6 Conclusion
The central finding of this work is that MTL’s value in materials science depends critically on the prediction goal. For high-precision property characterization—where accurate quantitative predictions guide material selection and design—independent single-task models remain superior. For high-throughput screening and discovery—where the goal is to identify promising candidates without missing potential hits—MTL offers meaningful benefits through improved recall, despite reduced precision. This distinction provides practitioners with an actionable framework for deciding when to apply MTL in their materials discovery pipelines.
Our systematic investigation of MTL for metal alloy property prediction, framed around three research questions, yields the following answers:
RQ1 (The Myth): Physical relatedness does not guarantee positive transfer. Resistivity and hardness, despite sharing atomic-level dependencies, follow fundamentally different functional forms (Nordheim-type polynomials vs. power-law lattice distortion), creating gradient directions misaligned by 66–75∘. Learned task relation graphs confirm this incompatibility with near-zero inter-task weights (0.006). The assumption that “properties sharing physics should share representations” is demonstrably false for this property combination.
RQ2 (The Stress Test): Extreme data imbalance (65:1) compounds functional mismatch to produce severe negative transfer. The majority task dominates the shared backbone, leaving minority tasks to fit their different functional forms using mismatched representations. This manifests as a 16.6% drop in hardness under standard MTL. However, gradient conflict mitigation techniques—particularly PCGrad (+12.4%) and LDS+GradNorm—can substantially recover performance, confirming that the imbalance is addressable with appropriate optimization strategies.
RQ3 (The Utility): Despite regression degradation, MTL provides practical utility for classification/screening tasks. The 17% improvement in amorphous-forming ability recall means fewer promising candidates are missed during high-throughput screening. The cross-entropy loss scales logarithmically and naturally bounds gradients, making classification robust to the same imbalance that destabilizes regression.
Based on these findings, we offer preliminary guidelines for practitioners working with similar datasets:
@@
Use independent models when precise property prediction is required for material characterization
@@
Use MTL for classification/screening tasks where improved recall is valuable for discovery
@@
Validate MTL assumptions empirically rather than assuming positive transfer from shared physics
@@
Apply PCGrad when maximizing minority-task performance is the priority
@@
Use LDS+GradNorm when balanced multi-task performance is desired
@@
We hypothesize—but do not validate—that organizing properties by physical mechanism (electronic, mechanical, thermodynamic) may predict MTL compatibility: within-mechanism property pairs may benefit from joint learning, while cross-mechanism combinations require explicit validation. However, our derived property experiment demonstrates that shared physics alone is insufficient—auxiliary tasks must provide informationally independent supervision. Testing this refined hypothesis with independently measured within-mechanism property pairs remains an important direction for future work.
We emphasize that our results are specific to the AI-Hub Korea metal alloy dataset and the three properties studied. Recent contradictory findings in the literature—where MTL improved rather than degraded performance on different material systems—underscore that our observations may not generalize broadly. We present this work as an empirical case study documenting conditions under which MTL can fail, rather than as universal principles. The key contribution is not the specific numerical results, but the demonstration that when to use MTL matters as much as how—a nuance often overlooked in materials informatics, where “more data” is frequently assumed to be universally beneficial.
@@
The metal alloy property datasets used in this study are available from the AI-Hub Korea platform [8]. Access requires free registration. The datasets include: (1) electrical resistivity measurements (52,388 samples), (2) Vickers hardness measurements (800 samples), and (3) amorphous-forming ability labels (840 samples).